# Online Resource Allocation via Static Bundle Pricing

### v1

@@ 1.1 Technical Overview and Contributions
In the $d$-single-minded setting, there is a set $\mathcal{M}$ of $m$ items, each with multiplicity specified by a capacity function $c:\mathcal{M}\to\mathbb{Z}_{>0}$. Buyers arrive online, each drawing a value $v_{b}\sim\mathcal{D}_{b}$, and derive utility $v_{b}$ if and only if they receive their entire desired bundle $S_{b}$, where $|S_{b}|\leq d$. Both the bundle $S_{b}$ and the distribution $\mathcal{D}_{b}$ are known in advance.
In this setting, we scale the item capacities in the LP relaxation by a factor $\gamma=e(10d)^{1/B}.$ Using a concentration argument, we show that for any fixed item, the probability that it becomes congested is at most $1/(10d)$. Taking a union bound over the at most $d$ items contained in a buyer’s desired bundle, we obtain that any bundle allocated by the unconstrained algorithm is congested with probability at most $0.1$. We leverage this bound to show that the expected value lost due to blocked buyers is at most a $1/10$ fraction of the value of the scaled fractional optimum. Combined with the properties of the unconstrained algorithm, this yields the following guarantee.
@@
For the $d$-single-minded case with minimum item multiplicity $B$, our mechanism achieves an $O(d^{1/B})$-competitive static and anonymous posted-price mechanism.
@@
We next consider the fully general single-minded setting, in which buyers may demand arbitrary bundles of items. Relative to the $d$-single-minded case, the absence of an a priori bound on bundle size requires more aggressive capacity scaling. Specifically, we scale down item capacities in the LP relaxation by a factor $\gamma=e(20m)^{1/(B+1)}.$ The slack created by this scaling allows us to show that the expected cardinality of the set of blocked buyers is at most $1/20$.
A further complication arises from the possibility of unbounded valuations: even a single blocked buyer with a very large value could dominate the welfare loss, rendering bounds on the number of blocked buyers insufficient. To address this issue, we augment the main mechanism with a simple small-market component that, with a carefully chosen probability, posts a single bundle at an extremely high price. This auxiliary mechanism captures a constant fraction of the welfare contributed by very large values, ensuring that heavy-tailed realizations cannot overwhelm the performance guarantee of the LP-guided menu. The main mechanism handles the structured, high-probability portion of the market, while the small-market mechanism isolates and stabilizes the contribution of extreme values. Together, the combined mechanism yields the following guarantee.
@@
For general single-minded combinatorial auctions with minimum multiplicity $B$, our mechanism achieves an $O\!\left(m^{1/(B+1)}\right)$-competitive static and anonymous posted-price mechanism.
Graph Setting.
Finally, we extend our framework to the network-routing domain, in which each buyer seeks to route a unit of flow from a source node $s$ to a target node $t$ in a capacitated graph $G=(V,E)$. We view this setting as a natural generalization of the general single-minded case: rather than demanding a single bundle, each buyer can be satisfied by any feasible $s$-$t$ path. The challenge arising by the extension is that the mechanism must select a specific routing for each buyer.
Once again, the structure of the LP relaxation plays a central role. We conceptually treat each source-target pair $(s,t)$ as an abstract bundle and group buyers according to their requested pair. Applying the same menu-construction procedure as in the general single-minded setting determines how many copies are posted to serve each $(s,t)$ pair, thereby inheriting the welfare guarantees of the unconstrained algorithm.
It remains to assign an explicit $s$-$t$ path to each such copy. To this end, we independently sample a path for each copy according to a carefully chosen distribution derived from the LP solution. Using a roulette-style sampling argument, we show that this randomized routing preserves the expected load guarantees on every edge, matching those obtained in the single-minded analysis. This yields the following result.
@@
For the graph-routing setting with minimum edge capacity $B$, our mechanism achieves an $O\!\left(m^{1/(B+1)}\right)$-competitive static posted-price mechanism, where $m=|E|$.
@@
Our impossibility results follow a standard lower-bound template for prophet inequalities (see, e.g., [chawla2009, babaioff]). In this template, buyers are partitioned into many groups of equal size $t$. Buyers arrive sequentially, each has value in $\{0,1\}$ with $\mathbb{P}\!\left[v=1\right]=1/t$, and feasibility is engineered so that we can select buyers from at most one group. Since the expected number of buyers realized in a fixed group is $1$, after committing to a group, the online algorithm can extract only constant expected value from that group and thus gains at most $O(1)$ value. In contrast, if we create about $t^{t}$ groups, then with constant probability there exists a fully active group (all $t$ buyers have value $1$), allowing the prophet to obtain welfare on the order of $t$. The main challenge is to embed this template into single-minded combinatorial auctions, where feasibility is induced by item capacities.
For unit capacities, we associate each group with a partition of the $m$ items into $t$ disjoint classes. The $t$ buyers in the group are single-minded, with buyer $i$ demanding exactly the $i$-th class. Disjointness within a partition guarantees that within a group all $t$ buyers can simultaneously be served. To prevent serving buyers from two different groups, we choose a family of partitions that are pairwise qualitatively independent (QI), a notion introduced by Rényi [renyi], where every class of one partition intersects every class of another. Then any allocation that serves one buyer from each of two groups must allocate some common item twice, contradicting unit capacity. This realizes the "at most one group" constraint of the template using item feasibility.
With $B$ copies per item, we generalize "at most one group" to "at most $B$ groups". The right combinatorial notion is (B+1)(B\!+\!1)-way qualitative independence: for any B+1B\!+\!1 partitions and any choice of one class from each, their intersection is nonempty. Using such a family, any allocation that serves buyers from B+1B\!+\!1 distinct groups must allocate some item to B+1B\!+\!1 buyers, contradicting the $B$-copy constraint; meanwhile, any fixed set of $B$ groups is simultaneously feasible because each item appears in at most one demanded class per group.
Poljak and Tuza [poljak] give the following lower bound on the number $r$-way qualitatively independent partitions of $m$ items into $t$ classes (denoted by $g(m,t,r)$):
$g(m,t,r)\geq\frac{er}{t}\cdot\exp\left(\frac{m}{rt^{r}}\right)$
thus, choosing
$t\;=\;\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor,$
via simple algebra we can prove that there exists a (B+1)(B\!+\!1)-way QI family of size at least $Bt^{t}$, thus we can instantiate $N=Bt^{t}$ groups. In our instance, each group is fully active with probability $\left(1/t\right)^{t}$ and if we let $X$ be the number of fully active groups we get that $\mathbb{E}\left[X\right]=N\cdot\left(1/t\right)^{t}=B$, and with constant probability it holds $X\geq B$. The prophet then serves all $tB$ active buyers, while any online algorithm is restricted to at most $B$ groups and extracts only constant expected value per served group.
Theorem 1.4.
For the general single-minded setting with $m$ items and minimum item capacity $B\geq 1$, every online algorithm has competitive ratio,
$\Omega\left(\left(\frac{m}{\ln m}\right)^{1/(B+2)}\right).$
To obtain a lower bound in the $d$-single-minded setting, we additionally require every demanded bundle to have size at most $d$. This leads us to introduce the notion of balanced QI partitions, in which every class has size $\leq d$. We show that the probabilistic construction of [poljak] extends to this balanced setting (sampling uniformly from balanced labelings), yielding large (B+1)(B\!+\!1)-way QI families with all classes of size about $m/t$. Setting $m=dt$ and choosing
$t\;=\;\left\lfloor\left(\frac{d}{2B\ln d}\right)^{1/(B+1)}\right\rfloor,$
the same group-commitment argument goes through.
Theorem 1.5.
For the $d$-single-minded setting with maximum desired bundle size $d$ and minimum item capacity $B\geq 1$, every online algorithm has competitive ratio,
$\Omega\left(\left(\frac{d}{\ln d}\right)^{1/(B+1)}\right).$
For unit capacities, there exist stronger combinatorial constructions that improve the dependence on $m$ beyond what the general probabilistic bound yields. Using the explicit construction of [poljak], we build $t^{t}$ pairwise QI $t$-partitions on $m=p^{3}$ items with $t=p$, for some sufficiently large prime $p$, eliminating the $\ln m$ loss.
Theorem 1.6.
For the single-minded setting with $m$ items and unit capacities, every online algorithm has competitive ratio,
$\Omega\left(m^{1/3}\right).$
Finally, we relate pairwise qualitative independence to the Alon-Alweiss measure (AAM) introduced by [AA20, SVW22], which yields even larger QI families on slightly fewer items and thus a mild exponent improvement.
Theorem 1.7.
For the single-minded setting with $m$ items and unit capacities, every online algorithm has competitive ratio,
$\Omega\left(m^{\,1/3+c/\log\log m}\right)$
for some absolute constant $c>0$.
1.2 Further Related Work
Online Contention Resolution Schemes.
We work with the ex-ante linear programming relaxation and interpret the fractional value assigned to each buyer-value pair as a target acceptance probability. The online realizations then play the role of a randomized rounding step. Our mechanism is responsible for enforcing feasibility ex-post, while preserving as much of the fractional optimum as possible.
This perspective is closely related to the framework of contention resolution schemes (CRSs), introduced by Chekuri, Vondrák, and Zenklusen for submodular maximization under combinatorial constraints [crs-chekuri]. In this framework, a CRS is given a fractional point $x^{*}$ in the relaxation polytope and samples a random set $R$ according to the product distribution with marginals $x^{*}$. Since the resulting set $R$ may violate feasibility constraints, the scheme subsequently discards selected elements to obtain a feasible subset $R^{\prime}\subseteq R$. Later, Feldman, Svensson, and Zenklusen [ocrs] introduced the online counterpart of this framework, termed online contention resolution schemes (OCRSs), which apply to a wide range of online settings, including prophet inequalities. In their work, they showed that the existence of a $c$-balanced OCRS implies a $c$-competitive prophet inequality. More recently, Lee and Singla [optimal-ocrs-singla] showed that a $c$-competitive prophet inequality also yields a $c$-competitive ex-ante OCRS.
Static versus Dynamic Pricing.
A related line of work is due to Rubinstein [rubinstein-beyond-matroids], who studied prophet inequalities under general downward-closed feasibility constraints. He obtained an $O(\log n\cdot\log r)$-competitive algorithm, where $n$ denotes the number of arriving elements (corresponding to buyers in our setting) and $r$ the maximum cardinality of any feasible solution. Rubinstein’s algorithm observes each realized value upon arrival and accepts the element if doing so does not decrease a carefully designed dynamic potential function. Since this potential function depends on the current accepted set, the resulting acceptance rule is inherently dynamic. Recently, Banihashem et al. [krysta] showed that any prophet inequality has an implementation using a (sequential) posted price mechanism.
Lower Bounds.
@@ 1.2 Further Related Work
The general $\Omega(\log n/\log\log n)$ lower bound in prophet inequalities with $n$ buyers that is used as a template for our main lower bound results has been introduced in other forms by Babaioff, Immorlica, and Kleinberg [babaioff] and by Chawla et al. [chawla2009]. For the $d$-single-minded cases, lower bounds for prophet inequalities under the intersection of $q$ partition matroids also imply hardness for $q$-single-minded prophet inequality instances: Kleinberg and Weinberg [klein-wein12] gave a $\Omega(\sqrt{q})$ lower bound and Saxena, Velusamy, and Weinberg [SVW22] strengthened this to $q^{1/2+\Omega(1/\log\log q)}$ using techniques of Alon and Alweiss [AA20].
2 Preliminaries
Model.
The input consists of a set of $m$ distinct items $\mathcal{M}$ and copies per item specified by the function $c:\mathcal{M}\to\mathbb{Z}_{>0}$. Let $B=\min_{e\in\mathcal{M}}c(e)$ denote the minimum number of copies among all items $e\in\mathcal{M}$. We are also given a set $\mathcal{N}$ of $n$ buyers. We index buyers by $b\in\{1,\ldots,n\}$. Each buyer $b$ is associated with a fixed set of items $S_{b}$ known to the seller. In addition, each buyer $b$ is associated with a valuation function $v^{(b)}:2^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ such that $v^{(b)}(S)=v_{b}$ if $S_{b}\subseteq S$ and $v^{(b)}(S)=0$ otherwise. Therefore, buyers are single-minded and thus fully described by the value $v_{b}$ and set $S_{b}$.
For the rest of the paper we will consider the Bayesian setting where $v_{b}$ is drawn from a distribution $\mathcal{D}_{b}$ that is known to the seller and independent across buyers. We will assume that the distributions are discrete with finite support $\mathcal{S}=\{0,\ldots,R\}$ with $R\in\mathbb{Z}_{>0}$, thus for each buyer $b$ and each $v\in\mathcal{S}$ we have $q_{b,v}>0$. In Appendix A we remove this assumption and we allow buyers to have different supports. Buyers arrive sequentially, each revealing their realized value upon arrival. We assume that the order of arrival is determined by an almighty adversary. Specifically, the almighty adversary can choose the arrival order after observing all realizations of the valuations and the random choices of the algorithm. Our objective is to maximize the social welfare, i.e., the sum of the values derived from the sets of items allocated to the buyers.
The class of mechanisms we consider is static and anonymous bundle pricing mechanisms. More specifically, before the arrival of the buyers, we offer a menu $M$ that contains multiple entries of the form $(T_{j},p_{j},c_{j})$, with $T_{j}\in 2^{\mathcal{M}}$, $p_{j}\in\mathbb{R}_{\geq 0}$ and $c_{j}\in\mathbb{Z}_{>0}$. In other words, we offer $c_{j}$ copies of the bundle of items $T_{j}$ at price $p_{j}$ each. Upon arrival, each buyer $b$ chooses a subset $J$ of available menu entries that maximizes their utility, where their utility is $v_{b}$ if $S_{b}\subseteq\bigcup_{j\in J}T_{j}$ and 0 otherwise, minus the total payment $\sum_{j\in J}p_{j}$. Note that a menu entry $(T_{j},p_{j},c_{j})$ can be chosen only if fewer than $c_{j}$ copies of this entry have been sold so far and copies of each item in $T_{j}$ are still available. The above class of mechanisms is static since it is determined upfront (before the realization of the buyers’ values) and does not change during the run of the algorithm. Also, it is anonymous because we offer the same menu to each buyer.
Notation.

### v2

@@ 2.1 General Framework for Static Bundle Pricing
In the $d$-single-minded setting, we assume that $|S_{b}|\leq d$ for all buyers. Scaling the item capacities in EA-LP by a factor $\gamma=e(10d)^{1/B}$ and using a concentration argument, we show that any bundle allocated by the unconstrained mechanism is blocked due to item capacities with probability at most $0.1$. Employing a careful analysis of the event that a buyer $b$ with realized value equal to the important value of their desired bundle $S_{b}$ is blocked, we show (in Lemma 7.2) that the expected value lost due to blocked buyers is at most a $\textnormal{{FracOpt}}_{\gamma}/10$ and obtain that:
@@
In the $d$-single-minded case with minimum item multiplicity $B$, our online static and anonymous bundle-pricing mechanism achieves a competitive ratio of $O(d^{1/B})$.
@@
We next consider general single-minded buyers that may demand bundles of arbitrary size. Using more aggressive capacity scaling, with $\gamma=e(20m)^{1/(B+1)}$, in EA-LP and concentration, we show that the expected number of blocked buyers is at most $1/20$.
A complication arises from the possibility of unbounded valuations: even a single blocked buyer with a very large value could dominate the welfare loss, rendering bounds on the number of blocked buyers insufficient. To address this issue, we augment the main mechanism with a simple small market component that, with a carefully chosen probability, posts a single bundle at price $2\textnormal{{FracOpt}}_{\gamma}$. This auxiliary part captures a constant fraction of the welfare contributed by very large values, ensuring that heavy-tailed realizations do not overwhelm the performance guarantee. The main mechanism deals with the structured, high-probability portion of the market, while the small-market component isolates and stabilizes the contribution of extreme values. Thus, we obtain:
@@
For general single-minded combinatorial auctions with minimum item multiplicity $B$, our online static and anonymous bundle-pricing mechanism is $O(m^{1/(B+1)})$-competitive.
Online Network Routing.
We extend our framework to online routing in a fixed directed network $G$. Instead of a bundle $S_{b}$, each buyer $b$ is now associated with two fixed known nodes, their source $s_{b}$ and their target $t_{b}$, and has value $v_{b}\sim\mathcal{D}_{b}$ for routing a unit of flow along any $s_{b}-t_{b}$ path.
For the menu construction, we conceptually treat each source-target pair $(s,t)$ as an abstract bundle and group buyers according to their requested pair. Applying the same menu construction as before, we determine how many copies are posted at each price for each $(s,t)$ pair. Thus, we directly inherit Property (1) of the unconstrained algorithm.
An additional challenge is that we should select a specific sts-t path for each allocated $(s,t)$-bundle. To this end, we independently sample a path for each $(s,t)$-bundle available according to a carefully chosen distribution, which is derived from the solution to EA-LP. Using a roulette-style sampling argument, we show that our randomized routing scheme over selected sts-t paths preserves the expected load guarantees on every edge and ensures Property (2) of the unconstrained algorithm. Relying on the approach towards proving Theorem 2.2, we obtain:
@@
For online routing in a network with $m$ edges, each with capacity at least $B$, our online static and anonymous bundle-pricing mechanism is $O(m^{1/(B+1)})$-competitive.
Incentive Compatibility.
An additional desirable property of our menu-based pricing scheme is price subadditivity. Namely, for every bundle $S$, the posted price of $S$ is no greater than the total price of any collection of bundles whose union includes $S$. This guarantees that a buyer who desires $S$ always prefers purchasing a copy of $S$ rather than assembling a larger collection of bundles covering $S$. Enforcing price subadditivity, however, is nontrivial. The prices suggested by EA-LP need not satisfy this property, and the supports of the buyer value distributions may not contain appropriate price levels that would allow for a subadditive menu without sacrificing performance. To address this issue, we first augment the support of each buyer’s value distribution with carefully chosen auxiliary price points. Then, before constructing our pricing menu, we apply a post-processing step to EA-LP solution that enforces subadditivity (the details can be found in Appendix B).
2.2 Lower Bounds
The proofs of our impossibility results follow a standard lower-bound template for prophet inequalities (see e.g., Chawla et al. (2010); Babaioff et al. (2007)). The buyers are partitioned into groups of equal size $t$. Buyers arrive sequentially, each has value $1$ with probability $1/t$ and 0 otherwise. Their bundles are engineered so that we can select buyers from at most one group. Since the expected number of buyers realized in a fixed group is $1$, after committing to a group, the online algorithm can extract an expected value of $O(1)$ from that group. In contrast, if we have about $t^{t}$ groups, with constant probability, there exists a group where all $t$ buyers have value $1$ (a.k.a. a fully active group), which results in an expected value of $\Theta(t)$ for the prophet. The main challenge is to embed the aforementioned bundle structure in the setting of single-minded combinatorial auctions, where feasibility is induced by item capacities.
For unit capacities, we associate each group with a partition of the $m$ items into $t$ disjoint classes. The $t$ buyers in the group are single-minded, with buyer $i$ demanding the $i$-th class. Disjointness within a partition guarantees that within a group, all $t$ buyers can simultaneously be accepted. To prevent accepting buyers from different groups, we choose a family of partitions that are pairwise qualitatively independent (QI), a notion introduced by Rényi (1970), where every class of one partition intersects every class of every other partition. Thus, no feasible allocation can accept buyers from different groups, thus realizing the “at most one group” constraint of the template.
With $B$ copies per item, we have the constraint of “at most $B$ groups”. The corresponding combinatorial notion is that of (B+1)(B\!+\!1)-way qualitative independence: for any B+1B\!+\!1 partitions and any choice of one class from each, their intersection is nonempty. Using such a family, any allocation that serves buyers from B+1B\!+\!1 distinct groups must violate the capacity constraint of some item. Meanwhile, any fixed set of $B$ groups is simultaneously feasible, because each item appears in at most one demanded class per group.
Poljak and Tuza (1989) gave a lower bound on the number of $r$-way qualitatively independent partitions of $m$ items into $t$ classes of $g(m,t,r)\geq\frac{r}{et}\cdot\exp\left({\frac{m}{rt^{r}}}\right)\,.$ Choosing $t=\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor$, we prove that there exists a (B+1)(B\!+\!1)-way QI family of size at least $Bt^{t}$; thus, we can instantiate $Bt^{t}$ groups of $t$ buyers each. In our construction, each group is fully active with probability $\left(1/t\right)^{t}$. Hence, the expected number of fully active groups is $B$, which implies that the expected social welfare of the prophet is $\Theta(tB)$. Due to our construction, any online algorithm is restricted to accepting buyers from at most $B$ groups and due to its online nature, can extract an expected value of $O(1)$ from each group. Therefore,
Theorem 2.4.
In the general single-minded setting with $m$ items, each with capacity at least $B$, the competitive ratio of any online algorithm is $\Omega((m/\ln m)^{1/(B+2)})$.
For unit capacities, we can slightly improve the lower bound of Theorem 2.4 first to $\Omega(m^{1/3})$, using Poljak’s explicit construction, and then to $m^{1/3+\Omega(1/\log\log m)}$, by relating pairwise qualitative independence to the Alon-Alweiss measure AAM introduced in Alon and Alweiss (2020); Saxena et al. (2023), which yields larger QI families on slightly fewer items.
To obtain a lower bound for the $d$-single-minded setting, we introduce the notion of balanced QI partitions, where every class has size at most $d$. We show that the probabilistic construction of Poljak and Tuza (1989) extends to the balanced setting, yielding large balanced (B+1)(B\!+\!1)-way QI families with all classes of size about $d=m/t$. Using $t=\left\lfloor(\frac{d}{2B\ln d})^{1/(B+1)}\right\rfloor$, we obtain that:
Theorem 2.5.
In the $d$-single-minded setting with $m$ items, each with capacity at least $B$, the competitive ratio of any online algorithm is $\Omega((d/\ln d)^{1/(B+1)})$.
3 Related Work
Online Contention Resolution Schemes.
Our general approach is inspired by the framework of contention resolution schemes (CRSs), introduced by Chekuri et al. (2011). A CRS is given a fractional point $x^{*}$ in the relaxation polytope and samples a random set $R$ according to the product distribution with marginals $x^{*}$. Since the resulting set $R$ may violate feasibility constraints, the CRS subsequently discards selected elements to obtain a feasible $R^{\prime}\subseteq R$. Subsequently, Feldman et al. (2016) introduced the online counterpart of CRSs, termed online contention resolution schemes (OCRSs), which apply to a wide range of online settings, including prophet inequalities. Feldman et al. (2016) showed that the existence of a $c$-balanced OCRS implies a $c$-competitive prophet inequality. More recently, Lee and Singla (2018) established the converse, i.e., that a $c$-competitive prophet inequality yields a $c$-competitive ex-ante OCRS.
A central concept in the OCRS framework is selectability. Informally, an OCRS is $c$-selectable with respect to a relaxation polytope $P$ if for every point $x\in P$, every active element is included in the final solution with probability at least $c$. However, restricting our attention to static bundle pricing mechanisms, selectability does not align well with anonymity. To illustrate, consider a single-item setting with two buyers, $a$ and $b$, each having deterministic value $1$. The relaxation polytope admits a feasible fractional solution that assigns mass $1/2$ to each buyer. Any static and anonymous pricing mechanism assigns a single price to the item, possibly drawn from a probability distribution. This price is accepted or rejected by both buyers. If the price is accepted, the adversary can always select buyer $a$ to receive the item, rendering meaningful selectability guarantees impossible.
Static versus Dynamic Pricing.
A related line of work is due to Rubinstein (2016), who studied prophet inequalities under general downward-closed feasibility constraints. He obtained an $O(\log n\cdot\log r)$-competitive algorithm, where $n$ denotes the number of arriving elements (corresponding to buyers in our setting) and $r$ the maximum cardinality of any feasible solution. Rubinstein’s algorithm observes each realized value upon arrival and accepts it, if doing so does not decrease a carefully designed dynamic potential function. Since this potential function depends on the currently accepted set, the resulting acceptance rule is inherently dynamic. Recently, Banihashem et al. (2024) proved that any prophet inequality can be implemented by a (sequential, non-static and non-anonymous) posted-price mechanism.
@@ 3 Related Work
The general $\Omega(\log n/\log\log n)$ lower bound in prophet inequalities with $n$ buyers, which is used as a template for the proofs of theorems 2.4 and 2.5, was introduced in other forms by Babaioff et al. (2007) and Chawla et al. (2010). Lower bounds for prophet inequalities under the intersection of $d$ partition matroids also imply lower bounds in our online $d$-single-minded setting. To this end, Kleinberg and Weinberg (2012) gave an $\Omega(\sqrt{d})$ lower bound and Saxena et al. (2023) strengthened this to $d^{1/2+\Omega(1/\log\log q)}$ using techniques from Alon and Alweiss (2020).
Computational (In)approximability.
Our mechanisms can be implemented in time polynomial in the number of buyers $n$ and the number of items $m$, as long as the maximum possible valuation is polynomially large in $n$ and $m$. Our approximation guarantees match (i) the computational complexity theoretic lower bounds of $\Omega(m^{1/(B+1)})$ for general single-minded combinatorial auctions Bartal et al. (2003) and $\Omega(\sqrt{m})$ for network routing in directed networks with unit edge capacities Guruswami et al. (2003); and (ii) the best known approximation guarantees of $O(d^{1/B})$ for $d$-single-minded combinatorial auctions Bansal et al. (2012) and $O(m^{1/(B+1)})$ for network routing in directed networks with edge capacity at least $B$ Kolliopoulos and Stein (2004).
4 Notation and Preliminaries
Our model is formally defined in Section 2. We next introduce some additional notation.
We let $k]\coloneqq\{1,\ldots,k\}$, for any $k\in\mathbb{Z}_{>0}$. We let $c(e)\in\mathbb{Z}_{>0}$ denote the multiplicity (a.k.a. capacity) of every item $e\in\mathcal{M}$. We assume that the buyers’ value distributions $\mathcal{D}_{b}$ are discrete with common finite support $\mathcal{U}=\{0,\ldots,R\}$ with $R\in\mathbb{Z}_{>0}$. Thus, for all $b\in\mathcal{N}$ and $v\in\mathcal{U}$, $q_{b,v}>0$. In Appendix A, we remove this assumption and allow buyers to have different supports.